Grade 9Trigonometry

sin3A÷1+2cos2A

oshi
11 Years agoGrade 9

3 Answers

Sharan
11 Years ago
sin3A           = 3sinA - 4sin3A            =    sinA(3 - 4sin2A)    =   sinA
1+2cos2A          1+2(1- 2sin2A)                 (3 - 4sin2A)
Sharan
11 Years ago
sin3A           = 3sinA - 4sin3A            =    sinA(3 - 4sin2A)    =   sinA
1+2cos2A          1+2(1- 2sin2A)                 (3 - 4sin2A)
PK NAIN
11 Years ago
 
sin3A           = 3sinA - 4sin3A            =    sinA(3 - 4sin2A)    =   sinA
1+2cos2A          1+2(1- 2sin2A)                 (3 - 4sin2A)